OpenStudy (anonymous):

How can you change an improper fraction of 3 21/20 to 4 1/20?

OpenStudy (anonymous):

you cant?

OpenStudy (anonymous):

This is what the book has, the original problem was 1 1/4 + 2 8/10 , which 1 1/4 = 1 5/10 and 2 8/10 = 2 16/20 so now the add 1 5/20 + 2 16/20 = 3 21/20 = 4 1/20 so i just want to know how and why it became 4 1/20.

OpenStudy (anonymous):

oooooooooh! sorryyyy i get it well what is 3 21/20? added together in one big fraction?

OpenStudy (anonymous):

its 81/20 wich simplifies into 4 and 1/20th:)